OpenStudy (anonymous):

use factoring to find the zeros of the function. use the zeros to find the vertex of its graph. y=x^2-x-132

Directrix (directrix):

x^2-x-132 = 0 (x-12)(x+11) = 0 If the product of two factors is 0, then either factor may be zero. x - 12 = 0 and x + 11 = 0 x = 12 and x = -11

OpenStudy (anonymous):

and vertex is half way between the zeros